# QueryCAD: Grounded Question Answering for CAD Models

### v1

@@ Abstract
©2024 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ I Introduction
In response to this, we propose QueryCAD, the first deep learning-based question-answering system for CAD models. QueryCAD is the first system to retrieve specific information like measurements, positions, or counts of features or parts of CAD models in response to natural-language questions. Given a free-text question, QueryCAD generates answers by grounding its reasoning process in structured engineering knowledge, directly retrieving measurements from the CAD model. Grounding natural-language queries in CAD data enables seamless integration of CAD models into deep learning-based applications and is particularly compatible with existing frameworks that build upon Large Language Models [2, 3]. QueryCAD is integrated into the MetaWizard robot program synthesis system [3] to enable the automatic generation of industrial robot programs given natural-language task specifications.
@@ II-C CAD Machining Feature Classification
All approaches share the common requirement of training a neural network for the classification task. This necessitates a large dataset, which is often generated synthetically [1, 31, 23] because existing large-scale real-world datasets [34, 35] lack the necessary labels [1]. While these methods are highly effective for machining classification and show a good generalization to new shapes, they are limited by their closed-set classification, making it difficult to adapt them for identifying additional feature types. Furthermore, these approaches typically operate at a low feature level, focusing on individual features rather than classifying higher-level features such as parts or more complex structures.
@@ III-A2 Instance Segmentation
To segment the rendered CAD model based on a free-text description, we utilize GroundingDINO [38] and Segment Anything Model (SAM) [39]. These models were selected due to the diversity of their training datasets and their demonstrated generalization capabilities across various domains [40, 41, 42]. In contrast, alternative approaches like Point-LLM [10] were trained on smaller, more specialized datasets, which limits their generalization potential. Additionally, our segmentation approach facilitates answering view-related questions by allowing us to render the model from specific angles.
GroundingDINO processes the rendered CAD image along with the free-text part description and outputs multiple bounding boxes with associated probabilities. We determined that a probability threshold of 30 % was optimal (see Ablation IV-B). Bounding boxes with probabilities below this threshold are discarded. Finally, we apply SAM [39] to each bounding box and rendered CAD image to generate precise segmentation masks. This combination of GroundingDINO and SAM was first proposed by Ren et al. [36].
GroundingDINO tends to select large portions or even the entire CAD model as a single bounding box. To address this, we filter out masks that cover more than 45 % of the rendered CAD model. The segmentation process can be seen in Figure 2.
@@ III-A3 Image to CAD Model Alignment
Figure 3: Pruning CAD faces to ensure only neighboring faces are selected as part. Starting with raw faces (2) derived from a 2D segmentation mask (1), we use ray casting to identify the face closest to the viewport (3). From this face, we select adjacent faces based on the adjacency graph (4) recursively. As a result, non-adjacent faces, particularly those further in the background, are discarded (5).
@@ IV-A Evaluation
Syntax addresses samples where the LLM generated invalid Python syntax, which was the case for none of the samples. Reasoning corresponds to samples where the proposed Python code features incorrect reasoning or hallucinations, which occurred once. If the Python code is valid, there are still two issues that can cause invalid answers: Masks, where incorrect parts were returned by SegCAD due to imprecise, missing or incorrect masks proposed by GroundedSAM or due to inaccuracies in the ray casting, which occurred 52 times. Lastly, CAD-Interface classifies answers where the CAD interface provided in the Python code was not powerful enough to answer the question, which was the case for 14 samples.

### v3

@@ I Introduction
In response to this, we propose QueryCAD, the first deep learning-based question-answering system for CAD models. QueryCAD is the first system to retrieve specific information like measurements, positions, or counts of features or parts of CAD models in response to natural-language questions. Given a free-text question, QueryCAD generates answers by grounding its reasoning process in structured engineering knowledge, directly retrieving measurements from the CAD model (see Fig. 1). Grounding natural-language queries in CAD data enables seamless integration of CAD models into deep learning-based applications and is particularly compatible with existing frameworks that build upon Large Language Models [2, 3]. QueryCAD is integrated into the MetaWizard robot program synthesis system [3] to enable the automatic generation of industrial robot programs given natural-language task specifications.
@@ II-C CAD Machining Feature Classification
All approaches share the common requirement of training a neural network for the classification task. This necessitates a large dataset, which is often generated synthetically [1, 31, 23] because existing large-scale real-world datasets [34, 35] lack the necessary labels [1]. While these methods are highly effective for machining classification and show a good generalization to new shapes, they are limited by their closed-set classification, making it difficult to adapt them for identifying additional feature types. Furthermore, these approaches typically operate at a low feature level, focusing on individual machining features rather than classifying higher-level features like parts or geometric structures.
@@ III-A2 Instance Segmentation
To segment the rendered CAD model based on a free-text description, we use GroundingDINO [38] and Segment Anything Model (SAM) [39], chosen for their diverse training datasets and strong generalization across domains [40, 41, 42]. In contrast, alternatives like Point-LLM [10] were trained on smaller, specialized datasets, limiting their generalization. Additionally, our approach enables answering view-related questions by rendering the model from specific angles. GroundingDINO processes the rendered CAD image and free-text part description, outputting multiple bounding boxes with associated probabilities. We found a 30 % probability threshold optimal (see Ablation IV-B). Bounding boxes with probabilities below this threshold are discarded. Finally, SAM [39] refines each bounding box and rendered CAD image into segmentation masks. This combination of GroundingDINO and SAM was first proposed by Ren et al. [36]. GroundingDINO tends to select large portions or even the entire CAD model as a single bounding box. To address this, we filter out masks that cover more than 45 % of the rendered CAD model. The segmentation process can be seen in Figure 2.
@@ III-A3 Image to CAD Model Alignment
Figure 3: Pruning CAD faces to retain only neighboring ones. From raw faces (2) derived from a 2D mask (1), ray casting identifies the closest face (3). Adjacent faces are then selected recursively via the adjacency graph (4), discarding non-adjacent ones, i.e. ones visible through holes (1). The final pruned part (5) excludes the faces on the right, as they are distant from the viewport (3) and not adjacent to the masked faces on the left.
@@ IV-A Evaluation
Syntax addresses samples where the LLM generated invalid Python syntax, which was the case for none of the samples. Reasoning corresponds to samples where the proposed Python code features incorrect reasoning or hallucinations, which occurred once. If the Python code is valid, there are still two issues that can cause invalid answers: Masks, where incorrect parts were returned by SegCAD due to imprecise, missing or incorrect masks proposed by GroundedSAM (see Fig. 5 top, left) or due to inaccuracies in the ray casting, which occurred 52 times. Lastly, CAD-Interface classifies answers where the CAD interface provided in the Python code was not powerful enough to answer the question, which was the case for 14 samples.